

********************************************************************
UARS WEEKLY STATUS REPORT - 10 SEP 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004099 has been reprocessed and archived at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004099.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 28 August 2004 (Orbit 70939, GMT Day 241) through
03 September 2004 (Orbit 71044, GMT Day 247).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 245/1018 71006 POWER OFF
The beta angle ranged between the angles of 38.6 degrees to 58.6
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 63 degrees on 06 September 2004 (DOY 250).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.57 to 1.58 degrees C. Additional
battery performance data for each battery on 27 August and 03
September 2004 are provided in Appendix D.
The clock error ranged from 10.3 to -11.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On 10 August 2004, during the 223/0121 Z event, VT-6 was commanded due
to the differential voltage being greater than 180 mV for three
consecutive orbits. The differential voltage returned to nominal
values after two orbits subsequent to the change. The spacecraft
remained in VT-6 until approximately 245/1600, at which time it was
commanded back to VT-5. All trending of the power subsystem during
this time looks nominal and the FOT will continue to use VT-6 should
the differential climb again.
A new release of the CMS code to allow for SMA events was delivered to
the backup CMS this week. This new release was built to fix the file
transfer problem encountered in the previous build. The preliminary
testing showed the file transfer issue to be resolved, but the build
still requires all other process testing. The goal for completion is
the end of September.
MEETINGS
The FOT attended a meeting with Paul Ondrus on Wednesday September 01,
2004 at 11AM. The meeting was in reference to the HRDI attitude error
that FDF is working on and the current status of the differential
voltage issue.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 25 September 2004 (DOY 269).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 03 September 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70991 244/1103 04-246 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71006 245/1018 04-247 HALOE POWER OFF (HI TBAT TEMP)
71008 245/1418 04-248 MPS VT L6 TO VT L5
71011 245/1858 04-248 MPS BATSOC SENT
71016 246/0216 04-249 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71024 246/1554 04-250 OBC/CLK ZERO EXEC STATUS REPORT
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 03 September 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 03 September 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70991 244/1120 CLK ADJ FR 0.0 to 20.6 MSEC/DAY
71016 246/0219 CLK ADJ FR 20.6 to 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 03 September 2004
27 August 2004 - GMT Day 240 Beta = 35.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.73/-3.99 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-050.4 +5.79/+4.83 +26.8/-14.4 26.4 1.02 19.7
3 +00.0/-100.8 +4.51/+3.87 +25.6/-14.0 26.4 1.02 19.1
03 September 2004 - GMT Day 247 Beta = 58.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.62 +00.0/+00.0 07.0 0.00 0.0
2 +22.4/-044.8 +3.87/+2.29 +20.8/-13.6 26.4 1.04 16.3
3 +11.2/-044.8 +3.23/+1.66 +20.8/-14.4 26.4 1.03 16.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
71008 245/1418 L6 -> L5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 03 September 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-059 FOT 71015 246/0136 N/A OBC/CLK adj cmd problem
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-050 FOT See List August 04 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-059 FOT 71015 246/0136 N/A OBC/CLK adj cmd problem